OpenStudy (pawanyadav):

Physics question :- Three uniform copper wires have their lengths in ratio 1:4:5 and their mass in ratio 5:2:1 their electrical resistance will be in the ratio..

OpenStudy (pawanyadav):

Please no qualified helper will help in this

OpenStudy (pawanyadav):

Sorry for thid

Parth (parthkohli):

OK, I won't.

OpenStudy (pawanyadav):

Thanks

Parth (parthkohli):

Why though? lol

OpenStudy (michele_laino):

hint: If such wires have the same cross sectional area, then their electrical resistance is proportional to the respectively length of such wires

OpenStudy (pawanyadav):

Because I want to know the views of my friends like me...

OpenStudy (michele_laino):

nevertheless that is not the case of your exercise

Parth (parthkohli):

But I'm in 11th grade.

Parth (parthkohli):

OK, let's get to it. We're given that they're made out of the same material. Their mass density is the same.\[\rho = \frac{m}{V}=\frac{m}{A \ell }\]\[\Rightarrow \rho_1 = \rho_2 = \rho_3 = \rho_{copper}\]\[\Rightarrow\frac{m_1}{A_1\ell_1}=\frac{m_2}{A_2\ell 2} = \frac{m_3}{A_3\ell_3 }\]Since you're given the relation between masses and lengths, you can obtain the ratio of cross-sectional areas. Then use \(R \propto \frac{\ell }{A}\)

OpenStudy (michele_laino):

we can express the electrical resistance \(R\) of a conductor, like a function of its mass \(m\) as below: \[\huge R = \frac{k}{{\delta \cdot {S^2}}}m\] where \(k\) is the resistivity, \(S\) is the cross sectional area, and \(\delta\) is the density of copper

OpenStudy (pawanyadav):

R~L/A Since resistivity of all the three wires are same as all are of same material So, where do we found the area?

OpenStudy (pawanyadav):

L/A

Parth (parthkohli):

We obtained the ratio of areas by using the fact that the volume-density (mass/volume) is the same across all wires. That's a vital assumption.

OpenStudy (pawanyadav):

Can you find me pls

OpenStudy (michele_laino):

being: \[\Large \frac{m}{{{S^2}}} = \frac{{{\delta ^2}{l^2}}}{m}\] after a simplification, we can write: \[\Large R = k\delta \frac{{{l^2}}}{m}\]

OpenStudy (pawanyadav):

OK done

OpenStudy (pawanyadav):

For L/A For all wire Multiplying numerator and denominator by L We get (L)(L)/(AL)=L^2/V Also the ratio of mass is the ratio of volume for every wire because density is constant (all wire are of copper) We need to find the ratio of (L)^2/M Now, 1/5:16/2:25/1 1/5:8:25 Now we are left with 5 in one of the denominator so multiply every ratio with 5 And we get 1×5/5=1 8×5=40 25×5=125 So the ratio of resistance is 1:40:125

Parth (parthkohli):

heyyyy that's what I said